# On the concentration distribution in turbulent thermals

### v1

@@ 5.2 PDFs of concentration
We have calculated the probability density function of the concentration as a function of time for all the experiments and simulations. The 100 bins of the PDF are logarithmically spaced between the minimal concentration values and 1 for the concentration and between $5\times 10^{-1}$ and ×$\times 10^{5}$ for the PDF of the concentration normalized by the mean concentration $\langle c\rangle$. Figure 7 represents these PDFs for a given experiment and a simulation, with the same $Re=3452$. The PDFs of $c$ for all others Reynolds are shown in Appendix B (see figures13, 14, 15). In figure 7, the first PDF that is plotted corresponds to $t/t_{g}\sim 10$ for the experiment, and $t/t_{g}=0$ for the simulation, for which the concentration field is available from the start. In both cases, the PDFs of concentration exhibit a drift toward smaller concentration values, in agreement with the time evolution of $\langle c\rangle$ and $\mathrm{max}(c)$ (see figure5a,b). For $t/t_{g}>10$, the shape of the PDF seems to be relatively constant with time and qualitatively consistent between the simulations and the experiments.
@@
Figure 7: PDF of $c$ as a function of $c$ and time $t/t_{g}$ for an experiment with $Re=3452$ (a,c) ($Pe=8.64\times 10^{6}$) and a simulation with the same $Re$ (b,d) ($Pe=3.4\times 10^{3}$). The top and bottom rows are log-log and semi-log plots, respectively.
@@
Figure 6 confirms the prediction of the dimensional analysis that the moments evolve as $\sim t^{-\frac{3}{2}k}$. Following the discussion of section 2.2, one should therefore expect the PDF of $c/\langle c\rangle$ to be independent of time. Figure 8 shows the PDFs of $c/\langle c\rangle$ in a log-log space and semi-log space for an experiment and a simulation, with $Re=3452$. Once normalized by $\langle c\rangle$, the PDFs of concentration for both simulation and experiment collapse (figure 8a,b), in agreement with the prediction of section 2.2.
@@
A similar behavior is observed for all Reynolds number (see figures 16, 17, 18). In the simulations, the PDF initially exhibits a different shape for the first few $t/t_{g}$ because the thermal is not yet fully developed. In addition, the concentration PDF remains nearly identical between simulations with the same Reynolds number but different resolutions (see in Appendix B figure 23).
5.3 Shape of the PDF of concentration
@@
Figure 8 shows that the PDF is linear for intermediate values of $c/\langle c\rangle$ when plotted in semi-log, suggesting that the PDF has the shape of a decreasing exponential, at least in this concentration range. However, care must be taken when characterizing the underlying PDF. Assuming it is of the form $p(X=c/\langle c\rangle)=\beta\exp{(-\lambda X)}$ as suggested by our data, the parameters $\lambda$ and $\beta$ are fixed by the requirement that the integral of $p(c/\langle c\rangle)$ over the concentration domain is 1, and that the expectancy (i.e. first moment) of $c/\langle c\rangle$ is 1. If $c\in[0,\infty[$, then $\lambda=\beta=1$ and, if exponential, the PDF of $c/\langle c\rangle$ is
@@
[…]
@@
The slope (in a semilogarithmic graph) of the apparent PDF is therefore dependent on the chosen threshold, and the effect is significant if the threshold $\xi$ is not $\ll 1$, which is the case here ($\xi=0.5$, see Section 4.2). One implication is that, for meaningful comparisons, PDFs from different experiments, simulations, and times must be calculated over the same domain of $c/\langle c\rangle(t)$ (i.e. same value of $\xi$).
As discussed in section 4.2, this threshold is set to $\xi=0.5$ at all times for all experiments and simulations. The cumulative distribution function associated with (24) is
@@
[…]
Comparison between the experimentally and numerically determined PDFs and the truncated exponential PDF given by (24) (grey line on figure 8) shows that the distribution of $c/\langle c\rangle$ is well described by the exponential distribution (24) for $c/\langle c\rangle$ smaller than 3 to 5, depending on $t$. According to equation (25), the probability of finding $c/\langle c\rangle<3$ in the thermal is equal to $99.3$%, which shows that the concentration in most of the thermal follows closely an exponential distribution, with the deviations at high concentrations being representative of less than 1% of the thermal cross-section surface area. Using the empirical cumulative density function calculated from the concentration maps rather than equation (25) gives similar results for experiments and simulations (see Appendix figures 20, 19, 21, 22).
Figure 8: PDF of $c/\langle c\rangle$ as a function of $c/\langle c\rangle$ and time $t/t_{g}$ for an experiment with $Re=3452$ (a, c) ($Pe=8.64\times 10^{6}$) and a simulation at the same $Re$ (b, d) ($Pe=3.4\times 10^{3}$). The top and bottom rows are log-log and semi-log plots, respectively. The gray lines show the exponential distribution defined by Eq. 24.
Once the PDFs of concentration have converged toward a largely self-similar distribution with time, we compare the time-averaged PDF and examine the effects of Reynolds or Péclet numbers (figure 9). The PDFs of $c/\langle c\rangle$ are averaged in time (for $t/t_{g}>10$) and averaged between runs for experiments (and between the two orthogonal plans for simulations). For intermediate values of concentration, the PDF slope seems to be independent of $Re$ or $Pe$ (figure 9a). For high values of concentration ($c/\langle c\rangle>5$, there are significant deviations from the exponential decrease, but without any clear trend with Reynolds or Péclet (figure 9b). These high values of concentration are related to the undiluted cores (see Section 4.2).
Figure 9: Temporal mean of the PDF of $c/\langle c\rangle$ as a function of $c/\langle c\rangle$ for all different $Re$ numbers for the experiments (solid lines) and the simulations (dashed lines). (a) and (b) are log-log and semi-log plots, respectively. The gray line is the truncated exponential distribution (Eq. (24)).
@@
In figure 6, the predicted moments are shown with grey dashed lines, using $c_{1}=2.0$, which is obtained by fitting experimental and numerical data with $\xi=0.5$. The good agreement between the prediction of Eq. 29 (which assumes an exponential PDF) and the data gives further support to our proposition that the concentration PDF is exponential. It also shows that the time evolution of all the moments of $c$, and therefore of the PDF, can indeed be accurately predicted with a single fitting parameter ($c_{1}$).
Figure 10a shows the moment of concentration $\langle c/\langle c\rangle)^{k}\rangle^{1/k}$ for different $k$ as a function of time. All moments are shifted along the y-axis for better readability. The moments $\langle c/\langle c\rangle)^{k}\rangle^{1/k}$ are almost constant with time in both experiments and simulations. Note that in figure 6, the evolution of the moment of concentration is well predicted by Eq. 29. Figure 10b shows the temporal mean of the moment of concentration $\overline{\langle(c/\langle c\rangle)^{k}\rangle}$ as a function of the predicted moment for a truncated exponential distribution (see Eq. 26). Moments of concentration are well explained by our truncated exponential distribution model, although the dispersion of the experimental and numerical values increases with $k$. Figure 10b shows no significant dependence on Reynolds or Schmidt numbers.
Figure 10: (a) Moment of concentration $\langle c/\langle c\rangle)^{k}\rangle^{1/k},\,\text{with}\,k=2,4,6,8$ as a function of time $t/t_{g}$ for the mean of three experiments at each $Re$ (solid lines) and the mean of two simulations at each $Re$ (dashed lines). The black lines show the theoretical $k$th moment $E[X^{k}]$ according to Eq. 26. Note that the moments are shifted along the y-axis to help the visualization. (b) Temporal mean of the moment of concentration $\overline{\langle c/\langle c\rangle)^{k}\rangle}$ against the predicted moment $\overline{\langle c/\langle c\rangle)^{k}\rangle}_{\mathrm{predicted}}$ (Eq. 26), with $k=2$ to $8$. The black stars denote the mean of the $k$th moment for all Reynolds (experiments and simulations).
5.4 On the effect of $Sc$
The evolution of the concentration moments in a turbulent thermal seems independent of the Schmidt number $Sc$ within the range $[1,2500]$ presented here. In addition, the statistical distribution of $c$ is also, at first order, independent of the diffusivity. However, differences do exist. First, we observe thicker structures or lamellae in the simulations than in the experiments (see figures 3 and  4). At any time, the diffusivity length scale is 50 times larger for the simulations than for the experiments.
Figure 9b reveals that the maximum concentration (compared to the mean concentration) is smaller in the $Sc=1$ simulations compared to the $Sc=2500$ experiments. This suggests that diffusion affects the evolution of the cores as maximum values of concentration correspond to the thermal cores (see figure 3c).
@@ 6 Discussion and conclusion
Early work on turbulent thermals - buoyant mass rising or sinking in a quiescent environment - primarily investigated their integral evolution, focusing on the concept of turbulent entrainment (Batchelor, 1954; Morton et al., 1956; Turner, 1986; Lecoanet & Jeevanjee, 2019). More recently, experimental studies examined the internal structure of thermals by measuring their density and vorticity content (Bond & Johari, 2005, 2010; Zhao et al., 2013). However, these studies have not examined the distributions of concentration in turbulent thermals, which is an insightful tool for investigating their internal structure and mixing dynamics.
@@
In this study, we performed laboratory experiments and numerical simulations of turbulent thermals under similar conditions, and have analyzed the evolution of the moments and distribution of concentration. Our results show that the concentration fields of turbulent thermals are self-similar in time, except at high concentrations. Far from the source, the concentration distribution, normalized by the average concentration, is independent of time in both experiments and simulations (figure 8). The normalized moments of concentration are consistently constant over time (figure 10a). As discussed in section 1, this aligns with the fact that both $Re$ and $Pe$ remain constant over time when defined with the instantaneous radius and velocity (Eq. 8), resulting in constant normalized concentration moments in the asymptotic regime (Eq. 21). Self-similarity also requires that the typical length scales of concentration heterogeneities remain at a fixed ratio of the thermal size, and therefore evolve in proportion to $r$. The natural length scales of the problem – thermal size, diffusion length, Batchelor scale $\ell_{B}=(D/\gamma)^{1/2}$, where $\gamma$ is the stretching rate of the concentration lamellae – appear to all have the same time dependence when far from the source, as shown below. The thermal radius increases as $t^{1/2}$ as predicted by equation 11. This is also the case for the diffusion length $(Dt)^{1/2}$. Concerning the Batchelor scale, dimensional analysis along the same lines as in Section 2.1 shows that, far from the source, the stretching rate should be of the form $\gamma=f_{\gamma}(Re)/t$, where $f_{\gamma}(Re)$ is an unknown function of $Re$. This predicts $\ell_{B}=f_{\gamma}(Re)^{-1/2}(Dt)^{1/2}$.
@@
The stationary nature of the distribution of $c/\langle c\rangle$, i.e. the self-similar behavior, can be qualitatively understood as resulting from a balance between the competing effects of homogenization within the thermal and continuous addition of ambient liquid through engulfment. This is well illustrated by the fact that the ratio of the standard deviation to the mean of concentration $\sqrt{\sigma^{2}}/\langle c\rangle$ is constant with time (equal to 1 for an exponential distribution), which contrasts with the evolution of stirred concentration fields in confined spaces, for which the ratio $\sqrt{\sigma^{2}}/\langle c\rangle$ tends towards 00 as mixing proceeds. However, the details of the mechanisms by which the stationary distribution is maintained have yet to be understood and modeled.
@@
Our interpretation of self-similarity contrasts with the findings of Zhao et al. (2013), who found that the overall thermal does not show a self-similar evolution, although the cores do evolve self-similarly. This conclusion relies on a different measure of self-similarity, where the stability of ensemble-averaged radial vorticity and density profiles is assessed over the development of the thermal. The profiles are normalized by their maximum value, located in the cores, which differs from our results, where we normalize the concentration distribution by the average concentration of the thermal. Since the maximum concentration decreases more slowly than the average concentration, this may explain why self-similarity is observed in the cores but not in the overall thermal. Normalizing the profiles by quantities averaged over the thermal could result in a self-similar evolution of the profiles, except in the core, which is consistent with our findings.
Our results reveal that the distribution of concentration is well approximated by a parameter-free exponential PDF (figure 9 and 10b), except at large concentrations where the distribution is relatively scattered. The departure from exponential behavior seems to be related to the high concentration cores of the thermal, and suggests a locally different mixing dynamics. The cores of the thermal are protected from mixing, whereas the peripheral region is mixed more efficiently (figure 2). This is consistent with the process of turbulent entrainment, which engulfs the surrounding fluid in the peripheral region toward the center of the thermal. The cores around which the thermal wraps remain relatively undiluted. This qualitative observation is supported by the concentration PDF, as the cores of the thermal correspond to the highest concentration values (figure 3 and 4). The slower decrease in maximum concentration compared to the average confirms that the cores are relatively protected from mixing (figure 5b). Furthermore, the maximum concentration decreases slower in the experiments than in the simulations (figure 5c), possibly indicating a diffusive mixing process since experiments have a lower diffusivity.
The distribution of concentration characterizes the fluid mixtures described as a superposition of concentration lamellae (Villermaux, 2019). In expanding systems, such as turbulent jets (Duplat et al., 2010), exponential tails result from the distribution of elongation events and mixing times among non-interacting lamellae. In confined systems, such as turbulent plumes (Villermaux & Duplat, 2003; Duplat & Villermaux, 2008) and porous media flow (Le Borgne et al., 2015), they arise from kinetic aggregation of lamellae, leading to the self-convolution of concentration distributions. In between these extreme situations, flows involving partial aggregation also produce exponential tails (Le Borgne et al., 2013, 2017). In the aggregation mechanism, the concentration PDF is related to the number of aggregations (number of lamellae in a given radius, see Duplat & Villermaux (2008); Le Borgne et al. (2017)), which increases with time as the mixing occurs. Since the PDF of concentration normalized by the mean is independent of time, our distribution may represent an effective aggregation number in the asymptotic regime. A dynamic equilibrium can be formed between a homogeneous fluid (high aggregation inside the thermals, where correlated aggregation might occur (Heyman et al., 2024)) and a fresh fluid (without aggregation) on a characteristic entrainment time scale. This specific configuration would lead to an apparent aggregation number close to 1 in the concentration distributions.
@@
An unexpected observation is the very weak effect of diffusion on the PDFs of concentration, as shown by the comparison of experiments and numerical simulations carried out at the same values of the Reynolds number, but differing by a factor of 2500 in terms of Peclet numbers. In both cases, the concentration distribution is well described by an exponential PDF. Across the range of Reynolds and Péclet numbers examined in our experiments and simulations, the concentration distributions are mostly independent of $Re$ and $Pe$ (figures 9 and 10), showing a weak effect of viscosity and diffusion on the distributions. At the same $Re$, the concentration maximum is higher in experiments ($Sc=2500$) than in numerical simulations ($Sc=1$), suggesting a marginal dependency on $Pe$, possibly through diffusive homogenization of the cores. However, diffusivity strongly affects the spatial structure of the concentration field, as shown by notably finer structures in the experiments than in simulations (figure 2c-d). This could be further analyzed with the distribution of the concentration increments and their moments, known as structure functions (Le Borgne et al., 2017; Villermaux, 2019). This approach would help characterize the mixing process that underlies the exponential distribution.
[…]
[Supplementary data] Here is a link to a video of the 4 experiments at each Reynolds. video
[…]

### v2

@@ V.2 PDFs of concentration
We have calculated the probability density function of the concentration as a function of time for all the experiments and simulations. The bins of the PDF are logarithmically spaced with a width in log-space of $0.066$ for the PDF of concentration $c/c_{0}$ and with a width in log-space of $0.12$ for the PDF of the concentration normalized by the mean concentration $\langle c\rangle$. Fig. 7 represents these PDFs for a given experiment and a simulation, with the same $Re=3452$. The PDFs of $c$ for all other Reynolds are shown in the Supplementary Materials [15]. In Fig. 7, the first PDF that is plotted corresponds to $t/t_{g}\sim 10$ for the experiment, and $t/t_{g}=0$ for the simulation, for which the concentration field is available from the start. In both cases, the PDFs of concentration exhibit a drift toward smaller concentration values, in agreement with the time evolution of $\langle c\rangle$ and $\mathrm{max}(c)$ (see figure 5a,b). For $t/t_{g}>10$, the shape of the PDF is to be relatively constant with time and qualitatively consistent between the simulations and the experiments.
@@
Figure 7: PDF of $c$ as a function of $c$ and time $t/t_{g}$ for an experiment with $Re=3452$ (a,c) ($Pe=8.64\times 10^{6}$) and a simulation with the same ReRe (b,d) ($Pe=3.4\times 10^{3}$). The top and bottom rows are log-log and semi-log plots, respectively.
@@
Fig. 6 confirms the prediction of the dimensional analysis that the moments evolve as $\sim t^{-\frac{3}{2}k}$. Following the discussion of section II.2, one should therefore expect the PDF of $c/\langle c\rangle$ to be independent of time. Fig. 8 shows the PDFs of $c/\langle c\rangle$ in a log-log space and semi-log space for an experiment and a simulation, with $Re=3452$. Once normalized by $\langle c\rangle$, the PDFs of concentration for both simulation and experiment collapse (Fig. 8a,b), in agreement with the prediction of section II.2.
@@
A similar behavior is observed for all Reynolds numbers (see Figs. 5-7 in Supplementary Material [15]). In the simulations, the PDF initially exhibits a different shape for the first few $t/t_{g}$ because the thermal is not yet fully developed. In addition, the concentration PDF remains nearly identical between simulations with the same Reynolds number but different resolutions (see Fig. 12 in Supplementary Material [15]).
V.3 Shape of the PDF of concentration
@@
Fig. 8 shows that the PDF has an exponential shape for intermediate values of $c/\langle c\rangle$ and for $(t/t_{g}>15)$ in both experiment and simulation cases. In the numerical simulations, the PDFs deviate significantly from the exponential at early times $(t/t_{g}<10)$, corresponding to the acceleration phase where self-similarity is not yet established. Similar deviations are expected in the experiments; however, they are not observed here as the experimental field of view was focused on the distal behavior far from the source, excluding the initial acceleration. Note that care must be taken when characterizing the underlying PDF. Assuming it is of the form $p(X=c/\langle c\rangle)=\beta\exp{(-\lambda X)}$ as suggested by our data, the parameters $\lambda$ and $\beta$ are fixed by the requirement that the integral of $p(c/\langle c\rangle)$ over the concentration domain is 1, and that the expectancy (i.e. first moment) of $c/\langle c\rangle$ is 1. If $c\in[0,\infty[$, then $\lambda=\beta=1$ and, if exponential, the PDF of $c/\langle c\rangle$ is
@@
However, the PDFs we have calculated are lower-bounded by the threshold value used to define the thermal, which is given by a fixed value $\xi$ of the ratio between the concentration and the mean. Imposing that $\int_{\xi}^{\infty}p(X=c/\langle c\rangle)dX=1$ and $\int_{\xi}^{\infty}p(X=c/\langle c\rangle)XdX=1$, we find that, if exponential, the truncated PDF of $c/\langle c\rangle$ is
@@
The slope (in a semilogarithmic graph) of the apparent PDF is therefore dependent on the chosen threshold, and the effect is significant if the threshold $\xi$ is not $\ll 1$, which is the case here ($\xi=0.5$, see Section IV.2). One implication is that, for meaningful comparisons, PDFs from different experiments, simulations, and times must be calculated over the same domain of $c/\langle c\rangle(t)$ (i.e. same value of $\xi$).
As discussed in section IV.2, this threshold is set to $\xi=0.5$ at all times for all experiments and simulations. The cumulative distribution function associated with (24) is
@@
$\Gamma(k+1,\xi)=\int_{\xi}^{\infty}x^{k}e^{-x}dx.$ (27)
Comparison between the experimentally and numerically determined PDF and the truncated exponential PDF given by (24) (gray line in Fig. 8) shows that the distribution of $c/\langle c\rangle$ is well described by the exponential distribution (24) for $c/\langle c\rangle$ smaller than 3 to 5, depending on $t$. According to equation (25), the probability of finding $c/\langle c\rangle<3$ in the thermal is equal to 99.399.3%, which shows that the concentration in most of the thermal follows closely an exponential distribution, with the deviations at high concentrations being representative of less than 1% of the thermal cross-section surface area. Using the empirical cumulative density function calculated from the concentration maps rather than equation (25) gives similar results for experiments and simulations (see Figs. 8-11 in Supplementary Materials [15]).
Figure 8: PDF of $c/\langle c\rangle$ as a function of $c/\langle c\rangle$ and time $t/t_{g}$ for an experiment with $Re=3452$ (a, c) ($Pe=8.64\times 10^{6}$) and a simulation at the same ReRe (b, d) ($Pe=3.4\times 10^{3}$). The top and bottom rows are linear-linear and semi-log plots, respectively. The gray lines show the exponential distribution defined by Eq. 24.
Once the PDFs of concentration have converged towards a self-similar distribution with time, we compare the time-averaged PDF and examine the effects of Reynolds or Péclet numbers (Fig 9). The PDFs of $c/\langle c\rangle$ are averaged in time (for $t/t_{g}>10$) and averaged between runs for experiments (and between the two orthogonal plans for simulations). For intermediate concentration values ($c/\langle c\rangle>2$), the PDF slope seems to be independent of $Re$ or $Pe$ (Fig 9a). For high concentration values ($c/\langle c\rangle>2$, there are significant deviations from the exponential decrease, but without any clear trend with Reynolds or Péclet (Fig. 9b). These high values of concentration are related to the undiluted cores (see Section IV.2).
Figure 9: PDF of time-averaged $c/\langle c\rangle$ as a function of $c/\langle c\rangle$ for all different ReRe numbers for the experiments (solid lines) and the simulations (dashed lines). (a) and (b) are linear-linear and semi-log plots, respectively. The gray line is the truncated exponential distribution (Eq. (24)).
@@
In Fig. 6, the predicted moments are shown with grey dashed lines, using $c_{1}=2.0$, which is obtained by fitting experimental and numerical data with $\xi=0.5$. The good agreement between the prediction of Eq. 29 (which assumes an exponential PDF) and the data gives further support to our proposition that the concentration PDF is exponential. It also shows that the time evolution of all the moments of $c$, and therefore of the PDF, can indeed be accurately predicted with a single fitting parameter ($c_{1}$).
Fig. 10a shows the moment of concentration $\langle c/\langle c\rangle)^{k}\rangle^{1/k}$ for different $k$ as a function of time. All moments are shifted along the y-axis for better readability. The moments $\langle c/\langle c\rangle)^{k}\rangle^{1/k}$ are almost constant with time in both experiments and simulations. Note that in Fig. 6, the evolution of the moment of concentration is well predicted by Eq. 29. Fig. 10b shows the temporal mean of the moment of concentration $\overline{\langle(c/\langle c\rangle)^{k}\rangle}$ as a function of the predicted moment for a truncated exponential distribution (see Eq. 26). Moments of concentration are well explained by our truncated exponential distribution model, although the dispersion of the experimental and numerical values increases with $k$. Fig. 10b shows no significant dependence on Reynolds or Schmidt numbers.
Figure 10: (a) Moment of concentration $\langle c/\langle c\rangle)^{k}\rangle^{1/k},\,\text{with}\,k=2,4,6,8$ as a function of time $t/t_{g}$ for the mean of three experiments at each ReRe (solid lines) and the mean of two simulations at each ReRe (dashed lines). The black lines show the theoretical $k$th moment $E[X^{k}]$ according to Eq. 26. Note that the moments are shifted along the y-axis to help the visualization. (b) Temporal mean of the moment of concentration $\overline{\langle c/\langle c\rangle)^{k}\rangle}$ against the predicted moment $\overline{\langle c/\langle c\rangle)^{k}\rangle}_{\mathrm{predicted}}$ (Eq. 26), with $k=2$ to $8$. The black stars denote the mean of the $k$th moment for all Reynolds (experiments and simulations).
V.4 On the effect of ScSc
The evolution of the concentration moments in a turbulent thermal is independent of the Schmidt number $Sc$ within the range $[1,2500]$ presented here. In addition, the statistical distribution of $c$ is also, at first order, independent of the diffusivity. However, differences do exist. First, we observe thicker structures or lamellae in the simulations than in the experiments (see Figs. 3 and  4). At any time, the diffusivity length scale is 50 times larger for the simulations than for the experiments.
Fig. 9b reveals that the maximum concentration (compared to the mean concentration) is smaller in the $Sc=1$ simulations compared to the $Sc=2500$ experiments. This suggests that diffusion affects the evolution of the cores as maximum values of concentration correspond to the thermal cores (see Fig. 3c).
@@ VI Discussion and conclusion
Early work on turbulent thermals - buoyant mass rising or sinking in a quiescent environment - primarily investigated their integral evolution, focusing on the concept of turbulent entrainment [2, 30, 38, 22]. More recently, experimental studies examined the internal structure of thermals by measuring their density and vorticity content [5, 6, 48]. However, these studies have not examined the distributions of concentration in turbulent thermals, which is an insightful tool for investigating their internal structure and mixing dynamics.
@@
In this study, we performed laboratory experiments and numerical simulations of turbulent thermals, and analyzed the evolution of the moments and distribution of concentration. Our results show that the concentration fields of turbulent thermals are self-similar in time, except at high concentrations. Far from the source, the concentration distribution, normalized by the average concentration, is independent of time in both experiments and simulations (Fig. 8). This is leading to normalized moments of concentration that are constant over time (Fig. 10a). As discussed in section I, this self-similar behavior aligns with the fact that both $Re$ and $Pe$ remain constant over time when defined with the instantaneous radius and velocity (Eq. 8), resulting in constant normalized concentration moments in the asymptotic regime (Eq. 21). Self-similarity also requires that the typical length scales of concentration heterogeneities remain at a fixed ratio of the thermal size, and therefore evolve in proportion to $r$. The natural length scales of the problem – thermal size, diffusion length, Batchelor scale $\ell_{B}=(D/\gamma)^{1/2}$, where $\gamma$ is the stretching rate of the concentration lamellae – appear to all have the same time dependence when far from the source, as shown below. The thermal radius increases as $t^{1/2}$ as predicted by equation 11. This is also the case for the diffusion length $(Dt)^{1/2}$. Dimensional analysis along the same lines as in Section II.1 shows that, far from the source, the stretching rate should be of the form $\gamma=f_{\gamma}(Re)/t$, where $f_{\gamma}(Re)$ is an unknown function of $Re$. Therefore, the Batchelor length scale $\ell_{B}$ evolves as $f_{\gamma}(Re)^{-1/2}(Dt)^{1/2}$.
@@
The stationary nature of the distribution of $c/\langle c\rangle$, i.e., the self-similar behavior, can be qualitatively understood as resulting from a balance between the competing effects of homogenization within the thermal and continuous addition of ambient liquid through engulfment. This is well illustrated by the fact that the ratio of the standard deviation to the mean of concentration $\sqrt{\sigma^{2}}/\langle c\rangle$ is constant with time (equal to 1 for an exponential distribution), which contrasts with the evolution of stirred concentration fields in confined spaces, for which the ratio $\sqrt{\sigma^{2}}/\langle c\rangle$ tends towards 0 as mixing proceeds. However, the details of the mechanisms by which the stationary distribution is maintained have yet to be understood and modeled.
@@
Our interpretation of self-similarity contrasts with the findings of Zhao et al. [48], who found that the overall thermal does not show a self-similar evolution, although the cores do evolve self-similarly. This conclusion relies on a different measure of self-similarity, where the stability of ensemble-averaged radial vorticity and density profiles is assessed over the development of the thermal. The profiles are normalized by their maximum value, located in the cores, which differs from our results, where we normalize the concentration distribution by the average concentration of the thermal. Since the maximum concentration decreases more slowly than the average concentration, this may explain why self-similarity is observed in the cores but not in the overall thermal. Normalizing the profiles by quantities averaged over the thermal could result in a self-similar evolution of the profiles, except in the core, which is consistent with our findings.
Our results show that the concentration distribution is well approximated by an exponential PDF (Fig. 9 and 10b), except at the largest concentrations where the distribution becomes more scattered. This exponential behavior is reminiscent of similar features observed in other flow configurations, such as the exponential tails in expanding systems (turbulent jets [10]) and in confined settings (turbulent plumes [40, 11] or porous-media flows [19]). In contrast, the PDF obtained here is stationary when rescaled by the time-dependent mean - a notable difference from these cases - owing to the continuous engulfment of fresh ambient fluid by the thermal. The concentration distribution characterizes the fluid mixing processes, which can be interpreted in terms of the evolution of concentration lamellae [41], governed by their stretching histories and by interactions between lamellae (whether they remain isolated or aggregate). In several related configurations, models describing lamellar stretching and interaction have been successfully developed [11, 21, 19, 20]. Identifying the mechanisms responsible for both the exponential shape of our PDF and its self-similarity will require further investigation to establish a quantitative model for the elongation dynamics.
We also observed a departure from exponential behavior, which is related to the high concentration cores of the thermal, and suggests locally different mixing dynamics. The cores of the thermal are protected from mixing, whereas the peripheral region is mixed more efficiently (Fig. 2). This separation of dynamics - where the core is ”protected” while the periphery is mixed — echoes the discussion by [34]. They highlight that in regions of high coherence, such as the vortex cores of our thermals, transport is likely governed by the conservation of vorticity, which effectively sequesters the scalar. In contrast, the peripheral regions follow a more ’Prandtl-like’ momentum exchange, where continuous engulfment facilitates rapid homogenization. Their discussion is consistent with our two observations: slower decrease in the maximum concentration compared to the maximum, and slower decrease in the experiments compared to the simulations. For our simulations with $Sc\sim 1$, concentration would be able to ”leak” outside the vortex cores due to the fatter concentration profile than the velocity profiles [34]. In contrast, for our experiments with $Sc>>1$, the concentration would be ”sequestered” inside the vortex cores [34].
@@
An unexpected observation is the very weak effect of diffusion on the PDFs of concentration, as shown by the comparison of experiments and numerical simulations carried out at the same values of the Reynolds number, but differing by a factor of 2500 in terms of Peclet numbers. In both cases, the concentration distribution is well described by the same exponential PDF. Across the range of Reynolds and Péclet numbers examined in our experiments and simulations, the concentration distributions are mostly independent of $Re$ and $Pe$ (Figs. 9 and 10), showing a weak effect of viscosity and diffusion on the distributions. This is consistent with the recent analysis by [34], who suggest that the difference between momentum ($u$) and scalar ($c$) transport is a mixing problem governed by the way eddies exchange mass with the environment. Their work demonstrates that once the flow reaches a regime where the ”smooth reservoir” of the environment is accessible via turbulent coarsening, the specific molecular diffusivity has a vanishingly small effect on the overall transport law. At the same $Re$, the concentration maximum is higher in experiments ($Sc=2500$) than in numerical simulations ($Sc=1$), suggesting a marginal dependency on $Pe$, possibly through diffusive homogenization of the cores. However, diffusivity strongly affects the spatial structure of the concentration field, as shown by notably finer structures in the experiments than in simulations (Fig. 2c-d). Further analyses of the distribution of the concentration increments and their moments, known as structure functions [20, 41], would help characterize the mixing process that underlies the exponential distribution.
Acknowledgements.